# ORCID: 0000-0002-0777-2295

### v1

@@ 4 Method
Our empirical study utilized an online experiment that simulated the Stack Overflow environment, followed by a post-survey. Participants were asked to rate answers based on their perceived helpfulness within a simulated Stack Overflow environment. The answer position and source code comment type (block comment, inline comment, no comment) were used as within-subjects factors in the experimental design. In addition to answer position, the experiment used the answer score as a between-subjects factor, with one group receiving the score treatment, displayed on the left side of the Stack Overflow answers, while the control group received the answers in the same order, but without scores. Participants were randomly assigned to either the treatment or control group. This experimental setup allowed us to isolate the effect of scores from content-based factors such as code comments, enabling a controlled comparison of perceived helpfulness across conditions. For each scenario, that is, for each Stack Overflow question, the authors developed three different code snippets, inspired by real Stack Overflow posts related to the question. Each answer consisted of a short explanation text and a source code snippet. We counterbalanced the comment types with source code snippets. Since each participant was presented with three code snippets for each question, this resulted in six possible combinations that were randomly selected to be presented to the participant for each scenario. To vary the order of the answers, we presented the three answers to the participant in random order.
@@ 5 Experimental Procedure and Post-survey
The online experiment was implemented using the questionnaire platform SoSci [^4] https://www.soscisurvey.de/ and was divided into three main parts: the introduction, the Stack Overflow simulation, and the post-survey. In the first part, the introduction, the participant was presented with a consent form that explicitly stated that participation was voluntary and anonymous, and that all data collected would be kept confidential. In addition, participants were given technical instructions, such as the requirement to use a PC or laptop (not a mobile device) and to ensure that the browser zoom was set to 100%. After the Stack Overflow simulation, the third part was a post-survey divided into three sections: (a) Stack Overflow questions, (b) general control questions related to expertise, and (c) demographic information. In the first part of the post-survey, participants were asked to rank Stack Overflow elements displayed as cards in order of importance when evaluating an answer. They used drag-and-drop to position each item on a scale from least to most important. The question was asked twice: First, we asked about the elements used in the experiment, that is, the answer score, blank code (i.e., code without comments), explanation text, inline source comments, and block source comments. By “explanation text,” we mean a description of the code placed at the beginning of an answer that is not formatted as inline or block comments. Second, we asked about the elements of Stack Overflow in general, which included the same elements as before, along two additional ones: the accepted answer marked as the best answer by the questioner and the answerer’s reputation score. We asked several questions about the user’s prior experience and expertise. Participants were asked about their active or passive experience with Stack Overflow, frequency of use, whether they had used a code snippet from the Q&A page, and their reasons for using Stack Overflow. They also provided information about their programming skills, including years of experience. Experience with different programming languages was rated on a 6-point Likert scale item from “never”, “rarely”, “seldom”, “sometimes”, “occasionally” to “most of the time.” An open text field was provided to list additional programming languages not included. In the last section of the questionnaire, participants provided demographic information such as gender, age, country of origin, education, and occupation or field of study.
@@ 6.1 Helpfulness of source code comment types (RQ1 and RQ2)
We used linear mixed models provided by the lme4555https://cran.r-project.org/web/packages/lme4/ and lmerTest [^6] https://cran.r-project.org/web/packages/lmerTest/ packages in R [^7] https://www.r-project.org to analyze the fixed factors: comment type (baseline = no comments, block comments, inline comments), expertise (experts vs. novices), position (first, second, third answer), and the presence of answer scores (no answer score vs. presence of answer score; first answer: score $10$, second answer: score $3$, and third answer: score 0). The dependent variable was the perceived helpfulness of an answer. We also examined interaction effects between comment type and expertise, as well as between position and user ratings.
@@ 8 Discussion and Implications
Since users’ assessment of the perceived helpfulness of code snippets can influence their decision to reuse them, our results have implications beyond Stack Overflow. Our study points to the context-dependence of perceived helpfulness and code comprehension, highlighting the need for further studies to understand how users decide to reuse code, whether from Stack Overflow answers or from outputs of AI-based assistants offering alternative solutions.
From a practical perspective, the findings suggest that developers and contributors on platforms such as Stack Overflow should prioritize adding code comments, with particular emphasis on block comments. These types of comments were perceived as more helpful by users, indicating their potential to increase the overall helpfulness and understandability of code snippets. The guidelines provided by Stack Overflow on writing good answers [^8] https://stackoverflow.com/help/how-to-answer do not mention code comments at all (as of July 2025). Our suggestion is to extend these instructions based on our findings.
Our study has implications beyond code reading and comprehension. With the widespread adoption of generative AI tools in software development (Anh Nguyen-Duc, 2024), the question arises of how to prompt these tools to generate the most helpful solutions for users’ needs. An interesting question is whether prompting generative AI tools to include block comments generally leads to the generation of more helpful code snippets than not prompting them to include this type of comment.
@@ 9 Conclusion
The goal of this study was to investigate the factors that influence how Stack Overflow users assess the helpfulness of code snippets and solutions. To this end, we developed a simulated Stack Overflow environment that closely resembled the platform’s layout and functionality. The primary focus was on the role of source code comments—specifically how different comment types (block, inline, or none) affect perceived helpfulness. In addition, we examined whether surface-level features inspired by the platform—namely answer position and score—affect users’ helpfulness assessment. We conducted an online experiment with 91 participants using this simulated environment. The results showed that code snippets with block comments were perceived as more helpful than those with inline comments, and both types were rated as more helpful than uncommented code. In contrast, answer order and scores had no significant effect on perceived helpfulness.

### v2

@@ 4 Method
Our empirical study utilized an online experiment that simulated the Stack Overflow environment, followed by a post-survey. Participants were asked to rate answers based on their perceived helpfulness within a simulated Stack Overflow environment. The answer position and source code comment type (block comment, inline comment, no comment) were used as within-subjects factors in the experimental design. In addition to answer position, the experiment used the answer score as a between-subjects factor, with one group receiving the score treatment, displayed on the left side of the Stack Overflow answers, while the control group received the answers in the same order, but without scores. Participants were randomly assigned to either the treatment or control group. This experimental setup allowed us to isolate the effect of scores from content-based factors such as code comments, enabling a controlled comparison of perceived helpfulness across conditions. For each scenario, that is, for each Stack Overflow question, the authors developed three different code snippets, inspired by real Stack Overflow posts related to the question. Each answer consisted of a short explanation text and a source code snippet. We counterbalanced the comment types with source code snippets. Since each participant was presented with three code snippets for each question, this resulted in six possible combinations that were randomly selected to be presented to a participant for each scenario. To vary the order of the answers, we presented the three answers to the participants in random order.
@@ 5 Experimental Procedure and Post-survey
The online experiment was implemented using the questionnaire platform SoSci [^4] https://www.soscisurvey.de/ and was divided into three main parts: the introduction, the Stack Overflow simulation, and the post-survey. In the first part, the introduction, the participant was presented with a consent form that explicitly stated that participation was voluntary and anonymous, and that all data collected would be kept confidential. In addition, participants were given technical instructions, such as the requirement to use a PC or laptop (not a mobile device) and to ensure that the browser zoom was set to 100%. After the Stack Overflow simulation, the third part was a post-survey divided into three sections: (a) Stack Overflow questions, (b) general control questions related to expertise, and (c) demographic information. In the first part of the post-survey, participants were asked to rank Stack Overflow elements displayed as cards in order of importance when evaluating an answer. They used drag-and-drop to position each item on a scale from least to most important. The question was asked twice: First, we asked about the elements used in the experiment, that is, the answer score, blank code (i.e., code without comments), explanation text, inline source comments, and block source comments. By “explanation text,” we mean a description of the code placed at the beginning of an answer that is not formatted as an inline or block comment. Second, we asked about the elements of Stack Overflow in general, which included the same elements as before, along two additional ones: the accepted answer marked as the best answer by the questioner and the answerer’s reputation score. We asked several questions about the user’s prior experience and expertise. Participants were asked about their active or passive experience with Stack Overflow, frequency of use, whether they had used a code snippet from the Q&A page, and their reasons for using Stack Overflow. They also provided information about their programming skills, including years of experience. Experience with different programming languages was rated on a 6-point Likert scale item from “never”, “rarely”, “seldom”, “sometimes”, “occasionally” to “most of the time.” An open text field was provided to list additional programming languages not included. In the last section of the questionnaire, participants provided demographic information such as gender, age, country of origin, education, and occupation or field of study.
@@ 6.1 Helpfulness of source code comment types (RQ1 and RQ2)
We used linear mixed models provided by the lme4555https://cran.r-project.org/web/packages/lme4/ and lmerTest [^6] https://cran.r-project.org/web/packages/lmerTest/ packages in R [^7] https://www.r-project.org to analyze the fixed factors: comment type (baseline = no comments, block comments, inline comments), expertise (experts vs. novices), position (first, second, third answer), and the presence of answer scores (no answer score vs. presence of answer score; first answer: score $10$, second answer: score $3$, and third answer: score 0). The dependent variable was the perceived helpfulness of an answer. We also examined interaction effects between comment type and expertise, as well as between position and score.
@@ 8 Discussion and Implications
Since users’ assessment of the perceived helpfulness of code snippets can influence their decision to reuse them, our results have implications beyond Stack Overflow. Our study points to the context-dependence of perceived helpfulness and code comprehension, highlighting the need for further studies to understand how users decide to reuse code—from Stack Overflow answers or from outputs of AI-based assistants offering alternative solutions.
From a practical perspective, the findings suggest that developers and contributors on platforms such as Stack Overflow should prioritize adding code comments, with a particular emphasis on block comments. These types of comments were perceived by users as more helpful, indicating their potential to increase the overall helpfulness and understandability of code snippets. The guidelines provided by Stack Overflow on writing good answers [^8] https://stackoverflow.com/help/how-to-answer do not mention code comments at all (as of July 2025). Our suggestion is to extend these instructions based on our findings.
Our study has implications beyond code reading and comprehension. With the widespread adoption of generative AI tools in software development (Anh Nguyen-Duc, 2024), the question arises of how to prompt these tools to generate the most helpful solutions matching users’ needs. An interesting question is whether prompting generative AI tools to include block comments generally leads to the generation of more helpful code snippets than not prompting them to include these comments.
@@ 9 Conclusion
The goal of this study was to investigate the factors that influence how Stack Overflow users assess the helpfulness of code snippets and solutions. To this end, we developed a simulated Stack Overflow environment that closely resembled the platform’s layout and functionality. The primary focus was on the role of source code comments—specifically how different comment types (block, inline, or none) affect perceived helpfulness. In addition, we examined whether surface-level features on the platform—namely answer position and score—affect users’ helpfulness assessment. We conducted an online experiment with 91 participants using this simulated environment. The results showed that code snippets with block comments were perceived as more helpful than those with inline comments, and both types were rated as more helpful than uncommented code. In contrast, answer order and scores had no significant effect on perceived helpfulness.